# Theories Without Models: Uncontrolled Idealizations in Particle Physics

### v1

@@ 3. Theoretical Models and Empirical Adequacy
According to van Frassen, in articulating a theory one begins with a set of fundamental principles and laws and, without taking into consideration any kind of empirical data, constructs a set of models which can be deductively derived from these principles. These models are then compared to the ‘appearances’, i.e., the structures which can be described in experimental and measurement reports. A physical theory is empirically adequate ‘if it has some model such that all appearances are isomorphic to empirical substructures of that model’ (van Fraassen 1980). Van Fraassen’s emphasis on the importance of theoretical models of a theory stems from his strong views in favour of the semantic view, according to which the structure of a theory is nothing but its class of mathematical models. For van Fraassen (1980, 1989), the presentation of a scientific theory consists of a description of a class of state-space types, where state spaces are understood as a generic collection of mathematical objects (such as vectors, functions, and numbers) denoting the physical state of a system. Theoretical models are understood in this context as mathematical structures that are deductively derivable from the basic principles of the theory, denoting sequences of states that form a trajectory in the state space over time. The connection of these models to the physical world is described by (van Fraassen 2008, pp. 168-170) in terms of an isomorphic relationship between the theoretical models and the surface models of an experiment, i.e. the idealized models of the raw data obtained from experiments. [^7] See Lutz (2014, 2021) for interesting work on generalizing van Fraassen’s notion of empirical adequacy. Together with the observable phenomena, theoretical models occupy the central stage of constructive empiricism, since these are ‘the two poles of scientific understanding, for the empiricist [..] The former are the target of scientific representation and the latter its vehicle’ (van Fraassen 2008, p. 238).
A possible way to understand theoretical models, according to van Fraassen, is to think of them as specifying:
  1. (i)
    a family $N$ of observables (representing physical quantities) each with a range of possible values
  2. (ii)
    a set S of possible states (i.e. the model’s state space)
  3. (iii)
    a stochastic response function $P_{s}^{n}$ for each $n$ in $N$ and $s$ in $S$, which is a probability measure on the range of $n$.
@@
The members of the set $N$ of observables are functions from the model’s state space to the real line, assigning values to various physical quantities that are typically represented by the parameters and variables of the model. The state space $S$ comprises the set of all possible configurations of the system’s variables, and can be thought of geometrically as an N-dimensional space whose dimensions are defined by the number $N$ of observables in the model. A given state $s$ is thus represented by a point in the state space whose coordinates stand for a possible combination of the values of the physical quantities of the system represented in the model. Finally, the stochastic response function $P_{s}^{n}$ is the model’s specification that a (direct or indirect) measurement of a given observable $n$ will yield a specific value $n_{i}$, when the state is $s_{i}$. The overall idea is that a theoretical model fits (or is ‘isomorphically embeddable to’) a surface model only if the theoretical model has some state $s$ such that the function $P_{s}^{n}$ contains the surface state $P$ of the surface model, relative to the given identification of the measurement setups as measurement of the physical magnitudes $n$ (ibid. p.169). A physical phenomenon is thus represented both ‘from below’ via the surface model and ‘from above’ via the theoretical model. The fitting of these two models establishes the link of the theory with the natural world and eventually its empirical adequacy. A characteristic example of a theoretical model in physics comes from the classical harmonic oscillator which satisfies the theory of classical mechanics. The dynamic behaviour of the system described in this model is completely determined by the equations of the background theory and is in full accordance with the theory’s fundamental principles. However, this purely theoretical model does not allow the direct comparison of the model with the corresponding surface models of data derived from experiments. Hence, the additional construction of empirically relevant models – based on these theoretical models – is often necessary, to compare the quantities predicted by the models with the outcomes of experimental measurements. Theoretical models thus provide the mathematical basis on which further empirically relevant models are built with respect to a corresponding target system. The aim is to obtain theoretical predictions for experimental observables via these new models, in order to test the empirical adequacy of the models and consequently of the background theory.
@@
As van Fraassen notes (ibid. p.311), the process of arriving at a working model from a background theory can take several forms, often involving the use of mathematical idealizations. These idealizations typically occur in cases where the theory provides theoretical models for which no analytic solutions can be found and thus, further modifications are required in order to make the models mathematically tractable. For instance, the (theoretical) model of the ideal pendulum in classical mechanics provides the pendulum’s equation of motion
@@
where $\theta$ is the angle between the string and the vertical, $g$ is the local gravitational acceleration and $l$ is the length of the string. Equation (1) specifies, in principle, the aforementioned stochastic response function according to which a measurement of $\theta$ will yield a specific value $\theta_{1}$ given a state $s_{1}$, and is a second order non-linear differential equation for which no exact analytic solution can be found. Hence, the introduction of a mathematical idealization is necessary in order to arrive at a mathematically tractable form of the empirically relevant model which will eventually enable the comparison to the relevant surface model.
@@
which has a well known exact solution. Insofar as there are good mathematical reasons to justify that (2) holds for small angles, the introduced idealization is controllable and the linearized version of the ideal pendulum is still a theoretical model, in that it is derivable from the theory modulo some controlled modifications.
In what follows, we draw on the modelling techniques of perturbative quantum field theory to argue that in addition to phenomenological models whose construction aims in accommodating empirical results, there is another important class of non-theoretical models in physics whose construction comes from the use of non-controllable idealizations aiming to make them mathematically tractable and empirically relevant. A careful analysis of the modelling methodology of quantum electrodynamics within perturbative quantum field theory indicates that the empirically relevant cut-off models produced by regularization and renormalization techniques are detached from their corresponding theoretical models and the background theory, due to the fact that the required justification for the truncation of the series and the employment of the renormalization group method has not been established in practice. This fact breaks down the derivational relationship between the empirically relevant models and their background theory, leading to problems both for our understanding of the theory-model relationship and the empirical adquacy of the theory– according to van Fraassen’s view and more widely.
@@ 4. Models in Perturbative QFT
Perturbative Quantum Field Theory is the standard theoretical framework for constructing models for scattering processes of subatomic particles in particle physics. Realistic models of interactions are constructed by performing scattering theory calculation using specific quantum field theory Lagrangians or Hamiltonians, such as those for QED and QCD. These calculations allow the construction of empirically relevant models that are capable of providing theoretical predictions for physical quantities that can then be compared directly to experimental observables, such as cross sections and decay widths. A first difficulty in this process arises from the fact that the construction of four-dimensional interacting models via scattering theory leads to intractable Hamiltonians since the introduction of interacting terms typically leads to non-linear equations of motion for which no exact solutions can be found analytically. In order to make the interacting models mathematically tractable, physicists often appeal to perturbation theory, in which the total Hamiltonian $H$ of the system in question is split into a free and an interacting part:
@@
where $H_{0}$ is the free Hamiltonian whose eigenstates can be calculated analytically and $gV$ represents the Hamiltonian of the interacting model with an interaction potential $V$ parameterised by a coupling parameter $g$. Given that the coupling strength of the theory is weak – as is the case in QED and in high-energy QCD – this mathematical idealization is controllable, since it can be justified by safely assuming that the solutions of the total Hamiltonian are close enough to the well-known solutions given by the free Hamiltonian $H_{0}$. One of the most important quantities to be obtained by applying perturbation theory is the S-matrix, the operator that maps the initial state of a physical system undergoing a scattering process to the final state. The importance of the S-matrix for particle phenomenology experiments boils down to the close relationship of its elements – also referred to as scattering amplitudes $\mathcal{M}$ – to the scattering cross sections that are (indirectly) measured in these experiments. The S-matrix thus contains some of the theoretical predictions of the empirically relevant model of the theory upon which the empirical adequacy of the model can be tested. However, when perturbative methods are applied to realistic quantum field theories like QED, the final result is an expression of the scattering amplitudes as a power series of the (bare) coupling parameter $g_{0}$ which has the following generic form
@@
where $A_{n}$ stands for a multiple integral over momentum $k$ at each $n^{th}$ order term of the series. The main idea is that for those interactions where the coupling constant is relatively small, the first factor ${g_{0}}^{n}$ vanishes as the series proceeds to higher orders, and thus the contribution of higher order terms decreases significantly. All that is needed to calculate the coefficients of the series for each order (and consequently the S-matrix elements) is, therefore, to evaluate a set of multiple integrals over momentum space. This is where two significant problems arise however. The first problem concerns the fact that the total sum of the series is sometimes either not known to converge or known not to converge, and thus a legitimate question arises as to whether the series is indeed a good approximation to the quantity in question. Moreover, the number and the complexity of each of these integrals increases rapidly for higher orders and thus, in practice, only the sum of the first few terms is evaluated and compared to experimental results, often with great success. The second complication stems from the fact that each individual term of the series often appears to diverge either at very high energies (as $k\rightarrow\infty$) or at very low energies (as $k\rightarrow 0$) leading to the so-called ultraviolet and infrared divergences respectively. This is the notorious ‘problem of infinities’ in quantum field theory and, as we shall see, the solution to this anomaly comes from various regularization and renormalization techniques. The crucial question is whether the mathematical methods employed at this stage can be rigorously justified, making the introduced idealizations controllable, as in the simple case of the ideal pendulum. If so, then one is warranted to maintain that the final products, i.e. the empirically relevant models that are eventually put to test by comparison with experimental results, are theoretical models by virtue of being derivable from the background theory via controllable idealizations. Here is where we encounter a problem however, since the mathematical treatment of these two complications indicates that the introduced idealizations are uncontrolled and possibly uncontrollable due to the lack of a rigorous mathematical proof for (i) the Borel summability of the series to establish strong asymptoticity and (ii) the calculation of the corresponding beta-functions in renormalization group methods to establish asymptotic safety. As a result, the strong relationship between the final models and the background theory required by the semantic view breaks down. Let us consider each of these problems in more detail.
@@ 5.1. Divergent Series
The first complication concerns the large order behaviour of the perturbative series as a whole, independently of the offending integrals in the individual terms. In addition to the fact that the integrals in each term of the series diverge, there is often no evidence that the total sum of the series converges, or, to make things worse, there is strong evidence that the sum of the series at all orders diverges. Given the infinite number of terms and the increasing intractability of the integrals as one proceeds in higher-order terms, what happens in practice is that only the first few terms of the series are usually calculated and compared to experimental results, often by using different techniques for calculating each term of the series. Nevertheless, in several cases the sum of the first few series turns out to be in very close agreement with the experimental results, which leads us to the question of how the empirical success of the truncated series should be explained. [^8] The theoretical prediction for the anomalous magnetic moment of the electron is probably the most famous example of this practice. The first term of the series was first calculated by Schwinger (1948) and found to be in close agreement with the then available experimental results. The current state of the art only allows the calculation of the first five terms which requires – amongst other things – an evaluation of 12672 Feynman diagrams in the tenth-order perturbation theory (Kinoshita 2014). Miller (2023) discusses at length the case of divergent series in perturbative quantum field theories and notes that a possible justification for the truncation of the series after the first few terms and the explanation of the empirical success of the approximating series can be given in terms of strong asymptoticity. [^9] As Miller notes, this explanation was also given by Dyson (1952) himself in his arguments for the divergence of the perturbative series. Unlike convergent series, when a divergent asymptotic series is used for approximating a function, the sum of the first few terms is very close to the exact value of the function, and as one includes more and more terms to the sum, the value of the series increasingly diverges until it becomes infinitely different. Hence, if the perturbative series for the matrix elements of a meson-meson scattering process is indeed asymptotic to the unknown exact solution, it should be no surprise that the sum of the first few terms is often found to be in agreement with experimental results. The problem, however, is that while a given function can only have one asymptotic expansion, the converse is not true. That is, an asymptotic expansion can correspond to multiple functions, and thus it is possible that two or more different functions have the same asymptotic expansion. In fact, it is possible for an infinite number of functions with an entirely different set of solutions to share the same asymptotic function. Hence, even if a perturbative expansion in QED is indeed asymptotic to an exact solution, it does not uniquely specify what that function is. The only possible way to uniquely determine a function via an asymptotic series is to show that the series satisfies a strong asymptotic condition, which, roughly speaking, requires the differences between the exact value of the function and its series representation to be appropriately small for every order of perturbation theory.
The condition of strong asymptoticity that shows that a perturbative expansion uniquely satisfies an exact solution is only established if and only if a series is Borel summable. Borel summation is a mathematical method that is particularly useful for summing divergent asymptotic series. Suppose a formal power series $\sum_{k=0}^{\infty}\alpha_{k}\lambda^{k}$ of a function $f(\lambda)$ and define the Borel transform $B$ to be its equivalent exponential series $B(\lambda t)\equiv\sum_{k=0}^{\infty}\frac{dk}{k!}t^{k}$. If the Borel transform can be used to produce a unique reconstruction of $f(\lambda)$ then the series is Borel summable. [^10] For a more rigorous discussion on Borel summation in the context of perturbative QFT see (Duncan 2012, pp. 400-3).
@@
The important point for our purposes is as follows: If one knows the asymptotic expansion of a function, the function can be uniquely reconstructed by Borell summation, thus showing that the series corresponds to a unique function determined by the background theory. Unfortunately, however, Borel summability has not only not been demonstrated but is in fact not expected to hold in phenomenologically interesting models of quantum field theory. Indeed, Duncan (2012) notes that ‘the property of Borel summability is an extremely fragile one, and one which we can hardly ever expect to be present in interesting relativistic field theories’ (p.403, emphasis added). For perturbative QED in particular, Borel summability is not expected to hold in general. [^11] Here we are following the discussion of Miller (2023). For further discussion in the physics literature see ’t Hooft (1979) Khuri (1981); Fischer (1997). Thus the explanation for the success of the series after the truncation of the first terms based on strong asymptoticity is only a conjecture. The relevant mathematical idealization is therefore uncontrolled and the derivational relationship between the empirically relevant models in perturbative QFT and the background theory is, as it stands, obstructed. We will return to this issue and consider the idea of exogenous control of the idealization via non-perturbative ‘instanton’ effects in Section 6.

### v2

@@ 3. Theoretical Models and Empirical Adequacy
According to van Fraassen, in articulating a theory one begins with a set of fundamental principles and laws and, without taking into consideration any kind of empirical data, constructs a set of models which can be deductively derived from these principles. These models are then compared to the ‘appearances’, i.e., the structures which can be described in experimental and measurement reports. A physical theory is empirically adequate ‘if it has some model such that all appearances are isomorphic to empirical substructures of that model’ (van Fraassen 1980). van Fraassen’s emphasis on the importance of theoretical models of a theory stems from his strong views in favour of the semantic view of the structure of scientific theories where the focus is on models, rather on the linguistic formulation of theories.999cf. van Fraassen (1989, p.217):‘[The semantic approach’s] difference from other approaches is largely one of attitude, orientation, and tactics rather than in doctrines or theses. The conviction involved is that concepts relating to models will be the more fruitful in the philosophical analysis of science’. For van Fraassen (1980, 1989), the presentation of a scientific theory consists of a description of a class of state-space types, where state spaces are understood as a generic collection of mathematical objects (such as vectors, functions, and numbers) denoting the physical state of a system. Theoretical models are understood in this context as mathematical structures that are deductively derivable from the basic principles of the theory, denoting sequences of states that form a trajectory in the state space over time. The connection of these models to the physical world is described by van Fraassen (2008, pp. 168-170) in terms of an isomorphic relationship between the theoretical models and the surface models of an experiment, i.e. the idealized models of the raw data obtained from experiments. [^10] See Lutz (2014, 2021) for interesting work on generalizing van Fraassen’s notion of empirical adequacy. Together with the observable phenomena, theoretical models occupy the central stage of constructive empiricism, since these are ‘the two poles of scientific understanding, for the empiricist [..] The former are the target of scientific representation and the latter its vehicle’ (van Fraassen 2008, p. 238). [^11] Things are of course not so straightforward when it comes to quantum and relativistic theories. For more details on van Fraassen’s treatment of these cases see van Fraassen (1980, Ch.6.3) and van Fraassen (1991, Ch.5).
@@
A characteristic example of a theoretical model in physics is the ideal pendulum in Newtonian mechanics. The dynamic behaviour of the system described in this model is completely determined by the equations of the background theory and is in full accordance with the theory’s fundamental principles. However, this purely theoretical model does not allow the (straightforward) comparison of the model with the corresponding surface models of data derived from experiments. Hence, the additional construction of empirically relevant models – based on these theoretical models – is often necessary, to compare the quantities predicted by the models with the outcomes of experimental measurements. Theoretical models thus provide the mathematical basis on which further empirically relevant models are built with respect to a corresponding target system. The aim is to obtain theoretical predictions for experimental observables via these new models, in order to test the empirical adequacy of the models and consequently of the background theory.
@@
As van Fraassen notes (ibid. p.311), the process of arriving at a working model from a background theory can take several forms, often involving the use of mathematical idealizations. These idealizations typically occur in cases where the theory provides theoretical models for which no analytic solutions can be found and thus, further modifications are required in order to make the models mathematically tractable. For instance, the (theoretical) model of the ideal pendulum in classical mechanics provides is given by the equation of motion
@@
where $\theta$ is the angle between the string and the vertical, $g$ is the local gravitational acceleration and $l$ is the length of the string. Equation (1) specifies, in principle, the aforementioned stochastic response function according to which a measurement of $\theta$ will yield a specific value $\theta_{1}$ given a state $s_{1}$, and is a second order non-linear differential equation for which exact analytic solution requires advanced methods of elliptic functions. However, the introduction of a mathematical idealization allows for a mathematically simple empirically relevant model which will eventually enable the comparison to the relevant surface model.
@@
which has a simple exact solution that can be derived via basic approaches to solving differential equations. Insofar as there are good mathematical reasons to justify that (2) holds for small angles, the introduced idealization is controllable and the linearized version of the ideal pendulum is still a theoretical model, in that it is derivable from the theory modulo some controlled modifications.
In what follows, we draw on the modelling techniques of perturbative quantum field theory to argue that in addition to phenomenological models whose construction aims in accommodating empirical results, there is another important class of non-theoretical models in physics whose construction comes from the use of non-controllable idealizations aiming to make them mathematically tractable and empirically relevant. A careful analysis of the modelling methodology of quantum electrodynamics within perturbative quantum field theory indicates that the empirically relevant cut-off models produced by regularization and renormalization techniques are detached from their corresponding theoretical models and the background theory, due to the fact that the required justification for the truncation of the series and the employment of the renormalization group method has not been established in practice. This fact breaks down the derivational relationship between the empirically relevant models and their background theory, leading to problems both for our understanding of the theory-model relationship and the empirical adequacy of the theory according to van Fraassen’s view and more widely.
@@ 4. Models in Perturbative QFT
Perturbative Quantum Field Theory is the standard theoretical framework for constructing models for scattering processes of subatomic particles in particle physics. Realistic models of interactions are constructed by performing scattering theory calculation using specific quantum field theory Lagrangians or Hamiltonians, such as those for QED and QCD. These calculations allow the construction of empirically relevant models that are capable of providing theoretical predictions for physical quantities that can then be compared directly to experimental observables, such as cross sections and decay widths. A first difficulty in this process arises from the fact that the construction of four-dimensional interacting models via scattering theory leads to intractable Hamiltonians since the introduction of interacting terms typically leads to non-linear equations of motion for which no exact solutions can be found analytically.
In order to make the interacting models mathematically tractable, physicists often appeal to perturbation theory, in which the total Hamiltonian $H$ of the system in question is assumed to be well-defined and equal to the sum of a free and an interacting part:
@@
where $H_{0}$ is the free Hamiltonian whose eigenstates can be calculated analytically and $gV$ represents the Hamiltonian of the interacting model with an interaction potential $V$ parameterised by a coupling parameter $g$. Given that the coupling strength of the theory is weak – as is the case in QED and in high-energy QCD – this mathematical idealization is controllable, since it can be justified by safely assuming that the solutions of the total Hamiltonian are close enough to the well-known solutions given by the free Hamiltonian $H_{0}$.
One of the most important quantities to be obtained by applying perturbation theory is the S-matrix, the operator that maps the initial state of a physical system undergoing a scattering process to the final state. The importance of the S-matrix for particle phenomenology experiments boils down to the close relationship of its elements – also referred to as scattering amplitudes $\mathcal{M}$ – to the scattering cross sections that are (indirectly) measured in these experiments. The S-matrix thus contains some of the theoretical predictions of the empirically relevant model of the theory upon which the empirical adequacy of the model can be tested.
When perturbative methods are applied to realistic quantum field theories like QED, the final result is an expression of the scattering amplitudes as a power series of the (bare) coupling parameter $g_{0}$ which has the following generic form
@@
where $A_{n}$ stands for a multiple integral over momentum $k$ at each $n^{th}$ order term of the series. For those interactions where the coupling constant is relatively small, the first factor ${g_{0}}^{n}$ vanishes as the series proceeds to higher orders, and thus the contribution of higher order terms decreases significantly. All that is needed to calculate the coefficients of the series for each order (and consequently the S-matrix elements) is, therefore, to evaluate a set of multiple integrals over momentum space.
This is where two significant problems arise. The first problem concerns the fact that the total sum of the series is sometimes either not known to converge or known not to converge, and thus a legitimate question arises as to whether the series is indeed a good approximation to the quantity in question. Moreover, the number and the complexity of each of these integrals increases rapidly for higher orders and thus, in practice, only the sum of the first few terms is evaluated and compared to experimental results, often with great success. The second complication stems from the fact that each individual term of the series often appears to diverge either at very high energies (as $k\rightarrow\infty$) or at very low energies (as $k\rightarrow 0$) leading to the so-called ultraviolet and infrared divergences respectively. This is the notorious ‘problem of infinities’ in quantum field theory and, as we shall see, the solution to this anomaly comes from various regularization and renormalization techniques.
The crucial question is whether the mathematical methods employed at this stage can be rigorously justified, making the introduced idealizations controllable, as in the simple case of the ideal pendulum. If so, then one is warranted to maintain that the final products, i.e. the empirically relevant models that are eventually put to test by comparison with experimental results, are theoretical models by virtue of being derivable from the background theory via controllable idealizations. Here is where we encounter a problem however, since the mathematical treatment of these two complications indicates that the introduced idealizations are uncontrolled and possibly uncontrollable due to the lack of a rigorous mathematical proof for (i) the Borel summability of the series to establish strong asymptoticity and (ii) the calculation of the corresponding beta-functions in renormalization group methods to establish asymptotic safety. As a result, the strong relationship between the final models and the background theory required by the semantic view breaks down. Let us consider each of these problems in more detail.
@@ 5.1. Divergent Series
The first complication concerns the large order behaviour of the perturbative series as a whole, independently of the offending integrals in the individual terms. In addition to the fact that the integrals in each term of the series diverge, there is often no evidence that the total sum of the series converges, or, to make things worse, there is strong evidence that the sum of the series at all orders diverges. Given the infinite number of terms and the increasing intractability of the integrals as one proceeds in higher-order terms, what happens in practice is that only the first few terms of the series are usually calculated and compared to experimental results, often by using different techniques for calculating each term of the series. Nevertheless, in several cases the sum of the first few series turns out to be in very close agreement with the experimental results, which leads us to the question of how the empirical success of the truncated series should be explained. [^12] The theoretical prediction for the anomalous magnetic moment of the electron is probably the most famous example of this practice. The first term of the series was first calculated by Schwinger (1948) and found to be in close agreement with the then available experimental results. The current state of the art only allows the calculation of the first five terms which requires – amongst other things – an evaluation of 12672 Feynman diagrams in the tenth-order perturbation theory (Kinoshita 2014).
Miller (2023) discusses at length the case of divergent series in perturbative quantum field theories and notes that a possible justification for the truncation of the series after the first few terms and the explanation of the empirical success of the approximating series can be given in terms of strong asymptoticity. [^13] As Miller notes, this explanation was also given by Dyson (1952) himself in his arguments for the divergence of the perturbative series. Unlike convergent series, when a divergent asymptotic series is used for approximating a function, the sum of the first few terms is very close to the exact value of the function, and as one includes more and more terms to the sum, the value of the series increasingly diverges until it becomes infinitely different. Hence, if the perturbative series for the matrix elements of a meson-meson scattering process is indeed asymptotic to the unknown exact solution, it should be no surprise that the sum of the first few terms is often found to be in agreement with experimental results.
The problem, however, is that while a given function can only have one asymptotic expansion, the converse is not true. That is, an asymptotic expansion can correspond to multiple functions, and thus it is possible that two or more different functions have the same asymptotic expansion. In fact, it is possible for an infinite number of functions with an entirely different set of solutions to share the same asymptotic function. Hence, even if a perturbative expansion in QED is indeed asymptotic to an exact solution, it does not uniquely specify what that function is. To uniquely determine a function via an asymptotic series one shows that the series satisfies a strong asymptotic condition, which, roughly speaking, requires the differences between the exact value of the function and its series representation to be appropriately small for every order of perturbation theory.
@@
The standard way to establish the condition of strong asymptoticity that shows that a perturbative expansion uniquely satisfies an exact solution is by showing that a series is Borel summable. Borel summation is a mathematical method that is particularly useful for summing divergent asymptotic series. Suppose a formal power series $\sum_{k=0}^{\infty}\alpha_{k}\lambda^{k}$ of a function $f(\lambda)$ and define the Borel transform $B$ to be its equivalent exponential series $B(\lambda t)\equiv\sum_{k=0}^{\infty}\frac{dk}{k!}t^{k}$. If the Borel transform can be used to produce a unique reconstruction of $f(\lambda)$ then the series is Borel summable. The important point for our purposes is as follows: If one knows the asymptotic expansion of a function, the function can be uniquely reconstructed by Borel summation, thus showing that the series corresponds to a unique function determined by the background theory. Unfortunately, however, Borel summability has not only not been demonstrated but is in fact not expected to hold in phenomenologically interesting models of quantum field theory. Indeed, Duncan (2012) notes that ‘the property of Borel summability is an extremely fragile one, and one which we can hardly ever expect to be present in interesting relativistic field theories’ (p.403, emphasis added). For perturbative QED in particular, Borel summability is not expected to hold in general. [^14] Here we are following the discussion of Miller (2023). For further discussion in the physics literature see ’t Hooft (1979) Khuri (1981); Fischer (1997).
It should be noted however, that Borel summation is not the only available reconstruction technique for connecting perturbative expansions with exact solutions. [^15] For a more rigorous discussion on Borel summation and possible alternatives in the context of perturbative QFT see (Duncan 2012, pp. 400-6). In fact, an alternative group of approaches under the name of ‘optimized perturbation theory’ retains some hope for potentially associating the asymptotic series with a unique function. An example of these methods is ‘linear $\delta$ expansion’ where the basic idea is to construct a series of approximates to the path integral which only require perturbative calculations, but nevertheless can be shown to converge rigorously to the exact answer. Roughly speaking, this technique can, in principle, establish that a sequence of approximations obtained by carrying out the $\delta$ expansion converges to an exact answer even if the original asymptotic expansion is not Borel summable. The problem however is that in the case of quantum field theories the radius of convergence is essentially unknown, and higher loop calculations are simply intractable so one is left to hope for a rapid convergence at low orders, usually less than five.
The upshot is that no available mathematical technique has thus far succeeded in associating the series with a unique solution. Thus the explanation for the success of the series after the truncation of the first terms based on strong asymptoticity is only a conjecture. The relevant mathematical idealization is therefore uncontrolled and the derivational relationship between the empirically relevant models in perturbative QFT and the background theory is, as it stands, obstructed. We will return to this issue and consider the idea of exogenous control of the idealization via non-perturbative ‘instanton’ effects in Section 6.